OpenStudy (anonymous):

Simplify 1/(1+(x/5)?

OpenStudy (anonymous):

\[\huge \frac{1}{1+\frac{x}{5}}\]?

OpenStudy (anonymous):

get rid of the compound fraction in one step by multiplying top and bottom by \(\huge 5\)

OpenStudy (anonymous):

Sorry, my computer froze. Am I supposed to multiply the entire fraction, or just x/5?

OpenStudy (anonymous):

all

OpenStudy (anonymous):

top and bottom

OpenStudy (anonymous):

like for example \[\frac{1}{3}=\frac{1}{3}\times \frac{5}{5}=\frac{5}{15}\]

OpenStudy (anonymous):

still have to be equal when you are done

OpenStudy (anonymous):

Oh, okay. I can solve it from here. Thank you!

OpenStudy (anonymous):

yw